OpenStudy (anonymous):

The function g(x) converts x grams to pounds, as shown. Which composite function expresses the number of pounds that can be lifted by a balloon containing x ft3 of air?

OpenStudy (anonymous):

g(x)=x/453.6

OpenStudy (anonymous):

When the air in a hot-air balloon is heated to 100 degrees Fahrenheit, each cubic foot (ft3) of air can lift about 7 grams of weight. Now that you know that air can lift weight, try answering the project questions. Answers a) g(f(x))=3175.2x b) g(f(x))= x/3175.2 c) g(f(x))= x/64.8